The spaceship uss Enterprise of mass 4.5 x 109 kg travelling with a velocity of 250 m/s is in a head on collision with a Klingon spaceship of mass 2.24 x 109 kg travelling at 120 m/s in the opposite direction. The collision results in the coupling of the two ships to form a wreckage. Calculate the common velocity of the wreckage immediately after the collision. 


1
Expert's answer
2021-05-19T17:03:32-0400
"m_Ev_E-m_Kv_K=(m_E+m_K)V\\
(4.5\cdot10^9)(250)-(2.24\cdot10^9)(120)\\=(4.5\cdot10^9+2.24\cdot10^9)V\\V=127\frac{m}{s}"